# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have demonstrated strong performance across various graph-based tasks by effectively capturing relational information between nodes. These models rely on iterative message passing to propagate node features, enabling nodes to aggregate information from their neighbors. Recent research has significantly improved the message-passing mechanism, enhancing GNN scalability on large-scale graphs. However, GNNs still face two main challenges: over-smoothing, where excessive message passing results in indistinguishable node representations, especially in deep networks incorporating high-order neighbors; and scalability issues, as traditional architectures suffer from high model complexity and increased inference time due to redundant information aggregation. This paper proposes a novel framework for large-scale graphs named ScaleGNN that simultaneously addresses both challenges by adaptively fusing multi-level graph features. We first construct neighbor matrices for each order, learning their relative information through trainable weights through an adaptive high-order feature fusion module. This allows the model to selectively emphasize informative high-order neighbors while reducing unnecessary computational costs. Additionally, we introduce a High-order redundant feature masking mechanism based on a Local Contribution Score (LCS), which enables the model to retain only the most relevant neighbors at each order, preventing redundant information propagation. Furthermore, low-order enhanced feature aggregation adaptively integrates low-order and high-order features based on task relevance, ensuring effective capture of both local and global structural information without excessive complexity. Extensive experiments on real-world datasets demonstrate that our approach consistently outperforms state-of-the-art GNN models in both accuracy and computational efficiency.
@@
Scalable Graph Neural Network, Over-smoothing, Large-scale graph
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable success in various graph-based learning tasks, such as node classification, link prediction, and recommender systems (Yu et al., 2022b, a; Chen et al., 2023; Li et al., 2023; Zhong et al., 2023). By leveraging the inherent structure of graphs, GNNs propagate information through iterative message passing, allowing nodes to aggregate features from their neighbors. This mechanism enables GNNs to effectively capture complex relationships and structural dependencies. However, despite their success, existing GNNs face two fundamental challenges when applied to large-scale graphs (Zhang et al., 2022a): over-smoothing and scalability issues.
@@ 1. Introduction
Over-smoothing occurs as the number of message-passing layers increases, causing node representations to become overly similar and lose their discriminative power. This issue is particularly severe when incorporating high-order neighbors (Jin et al., 2022; Zeng et al., 2023; He et al., 2023; Zhang et al., 2024; Gong et al., 2024), as their contributions may become redundant or even introduce noise, leading to performance degradation. Several existing techniques, such as residual connections, skip connections, and decoupling propagation from feature transformation, have been proposed to alleviate over-smoothing. However, these methods either introduce additional computational complexity or fail to effectively balance local and global information. Scalability (Zhang et al., 2018; Liu et al., 2020; Li et al., 2021; Wu et al., 2022; Ding et al., 2022; Zhao et al., 2023; Chen et al., 2024) is another major challenge, as many traditional GNN architectures struggle with high memory consumption and computational costs when applied to large-scale graphs. The primary reason is the exponential growth in the number of high-order neighbors, which leads to excessive information aggregation and redundant computations. Some recent works address scalability issues by adopting pre-computation-based techniques, mini-batch training, or sampling strategies. While these approaches improve efficiency, they often sacrifice performance due to incomplete neighborhood information.
@@
To overcome these challenges, we propose a novel framework towards scalable graph neural network via adapt high-order neighboring feature fusion named ScaleGNN, that fuses multi-hop graph features to address the over-smoothing issue. Specifically, we introduce a learnable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to adjust the relative importance of different neighborhood levels through trainable weight parameters. By regulating these weights, our approach selectively emphasizes the most informative high-order neighbors while minimizing the influence of less useful ones. Additionally, we introduce a task-aware feature fusion mechanism that adaptively integrates low-order and high-order features based on their relevance to the specific task. This ensures that our model effectively captures both local and global information without suffering from over-smoothing. Furthermore, we incorporate a selective neighbor filtering strategy based on a Local Contribution Score (LCS), which quantifies the relevance of high-order neighbors to the target node, ensuring that only the most useful neighbors contribute to feature aggregation. To validate the effectiveness of our approach, we conduct extensive experiments on multiple real-world graph datasets. The results demonstrate that our method consistently outperforms state-of-the-art GNN models in terms of both accuracy and efficiency, providing a scalable and robust solution to the over-smoothing problem in deep graph learning.
@@
  • We introduce a trainable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to selectively retain informative high-order neighbors while suppressing redundant or irrelevant ones. This ensures an effective balance between local and global information aggregation.
  • We propose a dynamic feature fusion mechanism that adaptively combines low-order and high-order features based on their relevance to the specific learning task, which allows the model to capture both local structure and long-range dependencies while mitigating over-smoothing.
  • We incorporate a selective high-order neighbor filtering strategy based on a local contribution score (LCS) to enhance computational efficiency, reducing unnecessary computational overhead by focusing only on the most valuable high-order neighbors
  • Extensive experiments on multiple real-world graph datasets demonstrate that our method consistently outperforms state-of-the-art GNN models in terms of both accuracy and efficiency, highlighting its scalability and practical applicability.
2. Related Work
2.1. Graph Neural Network
@@
Graph Neural Networks (GNNs) are specialized neural networks for graph deep learning. GCN (Kipf and Welling, 2016), an early model, performs message passing in each layer to aggregate 1-hop neighbor information, enriching vertex semantics. Stacking $K$ GCN layers allows GCNs to integrate information from neighbors within $K$ hops. GAT (Veličković et al., 2017) employs attention mechanisms to prioritize important neighbors during message passing. GraphSAGE (Hamilton et al., 2017) enhances scalability for large graphs with neighbor sampling for mini-batch training. SGC (Wu et al., 2019) simplifies multilayer GNNs by removing non-linearities and weight matrices between layers, leading to faster processing without sacrificing accuracy. APPNP (Gasteiger et al., 2018) introduces an improved propagation scheme inspired by personalized PageRank (Page, 1999), ensuring linear complexity via PageRank approximation. S2GC (Zhu and Koniusz, 2021) uses a modified Markov Diffusion Kernel for propagation, balancing global and local vertex contexts. However, these GNNs show limitations in many real-world graphs due to challenges like heterophily (Zhu et al., 2021; Pan and Kang, 2023; Li et al., 2024) and heterogeneity. In this paper, we focus on the challenge of heterogeneity, specifically that the aforementioned GNNs are tailored for homogeneous graphs and overlook the characteristics of heterogeneous graphs, where vertices and edges can be of different types and contribute differently. HGNNs have been developed to address this by accommodating the properties of heterogeneous graphs.
2.2. Heterogeneous Graph Neural Networks
@@
HGNNs can be classified into two categories: relation-wise and representation-wise styles. Relation-wise HGNNs identify neighbors based on different relations (meta-paths) and aggregate this information for vertex representations. HAN (Wang et al., 2019) uses hand-designed meta-paths and semantic attention to enhance GAT for neighbor aggregation. MAGNN (Fu et al., 2020) involves all vertices within a meta-path for better semantics. HetGNN (Zhang et al., 2019) samples neighbors via random walks and aggregates same-type neighbors with Bi-LSTMs. Representation-wise HGNNs perform multiple iterations of message passing and vertex updates. R-GCN (Schlichtkrull et al., 2018) extends GCN by using distinct transformation matrices for each 1-hop relation. RSHN (Zhu et al., 2019) uses a coarsened line graph for global edge-type embeddings. HetSANN (Hong et al., 2020) extends GAT with type-specific attention values. HGT (Hu et al., 2020) applies heterogeneous mutual attention with type-specific parameters. Simple-HGN (Lv et al., 2021) extends GAT by incorporating vertex features and learnable edge-type embeddings. HINormer (Mao et al., 2023) utilizes a Graph Transformer with a broader aggregation mechanism. These HGNNs are end-to-end models, but can be impractical for large graphs due to resource-intensive message passing during training.
Figure 1. The overview of the proposed ScaleGNN.
2.3. Scalable Graph Neural Networks
To address the efficiency challenges of HGNNs on large graphs, strategies like sampling (Ying et al., 2018; Jiang et al., 2021) and pre-computation-based methods have emerged. This paper focuses on pre-computation-based methods, which have shown superior performance and efficiency across datasets. One approach is building deep GNNs by decoupling feature propagation and non-linear transformation, enabling feature propagation without model parameter training. SGC (Wu et al., 2019)s eliminates nonlinearity between consecutive graph convolutional layers, improving scalability and efficiency. Following this, works like SIGN, S2GC, GBP, and GAMLP have been proposed to further enhance SGC’s performance while maintaining scalability. Another direction involves pre-computation-based scalable GNNs. NARS (Yu et al., 2020) mitigates high information loss by extracting multiple subgraphs based on sampled relation subsets and pre-computing on each subgraph rather than the entire graph. SeHGNN (Yang et al., 2023) exemplifies the relation-wise style by separately collecting neighbor information for target vertices across different relations within $K$ hops, reducing information loss but compromising efficiency. As the number of relations grows with $K$ hops, SeHGNN becomes inefficient, especially for $K>2$ on large graphs. RpHGNN (Hu et al., 2024) combines the low information loss of the relation-wise style with the efficiency of the representation-wise style.
3. Preliminary
@@
Generally, we consider an undirected large-scale graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\,\mathcal{X}\}$, where $\mathcal{V}$ is the collection of nodes, $\mathcal{E}$ is the collection of edges between the nodes, and $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the collection of node attributes, $f$ is the size of node feature vector. $C$ is the number of node classes, and $\mathbf{A}$ denotes the adjacency matrix of $\mathcal{G}$. $\mathbf{D}=diag(d_{1},d_{2},\ldots,d_{|\mathcal{V}|})\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the degree matrix of $\mathbf{A}$, where $d_{i}=\sum\nolimits_{{v_{j}}\in\mathcal{V}}{\mathbf{A}_{ij}}$ denotes the degree of node $v_{i}$. Given the labeled node set $\mathcal{V}_{l}$, our goal is to predict the labels for nodes in unlabeled set $\mathcal{V}_{u}=\mathcal{V}-\mathcal{V}_{l}$ with the supervision of $\mathcal{V}_{l}$.
4. Methodology

### v4

@@ Abstract
Graph Neural Networks (GNNs) have demonstrated impressive performance across diverse graph-based tasks by leveraging message passing to capture complex node relationships. However, when applied to large-scale real-world graphs, GNNs face two major challenges: First, it becomes increasingly difficult to ensure both scalability and efficiency, as the repeated aggregation of large neighborhoods leads to significant computational overhead; Second, the over-smoothing problem arises, where excessive or deep propagation makes node representations indistinguishable, severely hindering model expressiveness. To tackle these issues, we propose ScaleGNN, a novel framework that adaptively fuses multi-hop node features for both scalable and effective graph learning. First, we construct per-hop pure neighbor matrices that capture only the exclusive structural information at each hop, avoiding the redundancy of conventional aggregation. Then, an enhanced feature fusion strategy significantly balances low-order and high-order information, preserving both local detail and global correlations without incurring excessive complexity. To further reduce redundancy and over-smoothing, we introduce a Local Contribution Score (LCS)-based masking mechanism to filter out less relevant high-order neighbors, ensuring that only the most meaningful information is aggregated. In addition, learnable sparse constraints selectively integrate multi-hop valuable features, emphasizing the most informative high-order neighbors. Extensive experiments on real-world datasets demonstrate that ScaleGNN consistently outperforms state-of-the-art GNNs in both predictive accuracy and computational efficiency, highlighting its practical value for large-scale graph learning.
Index Terms:
@@
Scalable graph neural network, over-smoothing, large-scale graphs.
I Introduction
Graph Neural Networks (GNNs) have become the cornerstone of modern graph representation learning [1, 2], driving advances in applications such as node classification [3], link prediction [4, 5, 6], traffic prediction [7], and recommender systems [8, 9, 10]. The essence of GNNs lies in their iterative message-passing schemes [11, 12], whereby nodes update their representations by aggregating information from their local neighborhoods. This capability enables GNNs to naturally encode node features and capture node correlations.
As digital infrastructures evolve, graphs with millions or even billions of nodes and edges have become commonplace, magnifying the necessity for GNNs that are not only expressive but also computationally scalable [13, 14, 15, 16, 17]. In response, recent research communities have developed a rich array of scalable and deep GNN architectures, with significant innovations focused on balancing model expressiveness, computational efficiency, and training feasibility. One line of research is sampling-based scalable GNN methods such as GraphSAGE [18] and Cluster-GCN [19], which attempt to control computational costs by selecting subsets of neighbors during training, thereby reducing memory overhead and enabling mini-batch processing. These approaches help alleviate the burden of large-scale aggregation but may lose important structural information due to incomplete sampling. Another stream is pre-computation-based scalable approaches (including SGC [20], SIGN [21], and GAMLP [22]), which decouple neighborhood aggregation from parameter learning, allowing feature propagation to be computed in advance and greatly accelerating both training and inference[13, 14]. Meanwhile, a growing body of research has focused on deep and high-order GNNs (e.g., S2GC [23], GBP [24], RpHGNN [25]). They integrate information from multi-hop or high-order neighbors, thereby enriching the receptive field of each node and improving the capture of long-range dependencies. These advances have substantially expanded the applicability of GNNs to web-scale scenarios such as large-scale heterogeneous information networks [15, 16, 17].
@@ I Introduction
Despite these substantial developments, GNNs still face two persistent and intertwined challenges when deployed on large-scale and complex graphs. First, as the number of message-passing layers increases, GNNs suffer from the over-smoothing problem that node representations become overly similar and lose their discriminative power. This issue is particularly severe when incorporating high-order neighbors [26, 27, 28, 29, 30], as their contributions may become redundant or even introduce noise, leading to performance degradation. Although several techniques, such as residual connections [31], skip connections [20, 32], and decoupling propagation [33] from feature transformation, have been proposed to alleviate over-smoothing, they often introduce additional computational complexity or fail to effectively balance local and global information. Second, many traditional GNN architectures struggle with the scalability issue that models face high consumption and computational costs when applied to large-scale graphs [34, 35, 36, 37, 38, 39, 40]. The primary reason is the exponential growth in the number of high-order neighbors, which leads to excessive information aggregation and redundant computations. Some recent works address scalability issues [25] by adopting pre-computation-based techniques, mini-batch training, or sampling strategies. While these approaches improve efficiency, they often sacrifice performance due to incomplete multi-hop neighborhood information. Recently, plug-in modules such as RMask [41] have been introduced to further enhance scalable GNNs. These modules can be seamlessly integrated into frameworks like S2GC and GAMLP, and employ techniques such as noise masking or random walk-based neighbor selection to reduce redundancy and over-smoothing. While plug-in enhancements offer practical improvements, most still fundamentally inherit the aggregation schemes of their base models [26, 27, 28, 29, 30] and often lack adaptive, effective feature fusion of information from different neighborhood hops. Overall, despite advances in sampling-based, pre-computation-based, deep and high-order aggregation, and plug-in modules, scalability and over-smoothing issues still hinder scalable GNNs for large-scale and complex graphs. A unified solution resolving both over-smoothing and scalability remains essential for accurate and efficient large-scale graph learning.
@@
To overcome these challenges, we propose a novel framework towards scalable graph neural networks via adapting high-order neighboring feature fusion named ScaleGNN, which fuses multi-hop graph features to address the over-smoothing issue. Specifically, we introduce a learnable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to adjust the relative importance of different neighborhood orders through trainable weight parameters. By regulating these weights, our approach selectively emphasizes the most informative high-order neighbors while minimizing the influence of less useful ones. Additionally, we introduce a low-order enhanced feature fusion mechanism that adaptively integrates low-order and high-order features based on their relevance to the node classification task. This ensures that our ScaleGNN effectively captures both local and global information, mitigating the over-smoothing problem. Furthermore, we incorporate a selective neighbor masking strategy named Local Contribution Score (LCS) and impose LCS-based sparse constraints to prune less important high-order neighbors, which quantify the relevance of high-order neighbors to target nodes, ensuring that only the most useful neighboring nodes contribute to feature aggregation. To validate the effectiveness of ScaleGNN, we conduct extensive experiments on multiple real-world graph datasets. The results demonstrate that our ScaleGNN consistently outperforms state-of-the-art GNN models in terms of both accuracy and efficiency, providing a scalable and robust solution to scalability and over-smoothing problems in large-scale and complex graph learning.
@@
  • We design a trainable mechanism to construct and refine multi-hop neighbor matrices, effectively retaining informative high-order neighbors while filtering out redundancy and noise for better high-order information integration.
  • We propose a low-order enhanced fusion strategy that adaptively weights low-order and high-order features, enabling the model to capture both local detail and global correlations while alleviating over-smoothing.
  • We introduce the selective masking strategy named Local Contribution Score (LCS) to reduce multi-hop redundancy and impose sparse constraints based on LCS to prune unimportant high-order neighbors, significantly reducing computational cost with minimal performance loss.
  • Experiments on real-world graphs show that our model consistently outperforms state-of-the-art GNNs both in accuracy and efficiency, demonstrating superior scalability.
II Related Work
II-A Graph Neural Network
@@
Graph Neural Networks (GNNs) are specialized neural networks for graph deep learning. GCN [42], an early model, performs message passing in each layer to aggregate 1-hop neighbor information, enriching vertex semantics. Stacking $K$ GCN layers allows GCNs to integrate information from neighbors within $K$ hops. Graph Attention Network (GAT) [43] employs attention mechanisms to prioritize important neighbors during message passing. GraphSAGE [18] enhances scalability for large graphs with neighbor sampling for mini-batch training. SGC [20] simplifies multilayer GNNs by removing nonlinearities and weight matrices between layers, leading to faster processing without sacrificing accuracy. APPNP [44] introduces an improved propagation scheme inspired by personalized PageRank [45], ensuring linear complexity via PageRank approximation. S2GC [23] uses a modified Markov Diffusion Kernel for propagation, balancing global and local vertex contexts. However, these GNNs show limitations in many real-world graphs due to challenges like heterophily [46, 47, 48] and heterogeneity. In this paper, we focus on the challenge of heterogeneity, specifically that the aforementioned GNNs are tailored for homogeneous graphs and overlook the characteristics of heterogeneous graphs [49], where vertices and edges can be of different types and contribute differently. Heterogeneous Graph Neural Networks (HGNNs) have been developed to address this by accommodating the properties of heterogeneous graphs.
@@
HGNNs can be classified into two categories: relation-wise and representation-wise styles. Relation-wise HGNNs identify neighbors based on different relations (meta-paths) and aggregate this information for vertex representations. HAN [50] uses hand-designed meta-paths and semantic attention to enhance GAT for neighbor aggregation. MAGNN [51] involves all vertices within a meta-path for better semantics. HetGNN [52] samples neighbors via random walks and aggregates same-type neighbors with Bi-LSTMs. Representation-wise HGNNs perform multiple iterations of message passing and vertex updates. R-GCN [53] extends GCN by using distinct transformation matrices for each 1-hop relation. RSHN [54] uses a coarsened line graph for global edge-type embeddings. HetSANN [55] extends GAT with type-specific attention values. HGT [56] applies heterogeneous mutual attention with type-specific parameters. Simple-HGN [57] extends GAT by incorporating vertex features and learnable edge-type embeddings. HINormer [58] utilizes a Graph Transformer with a broader aggregation mechanism. HGAMLP [59] enhances heterogeneous graph learning by adaptively extracting and fusing local and global knowledge, while removing redundancy from high-order meta-paths, achieving state-of-the-art accuracy and efficiency on large-scale heterogeneous graphs. These HGNNs are end-to-end models but can be impractical for large graphs due to resource-intensive message passing during training.
Figure 1: The overview of the proposed ScaleGNN.
II-B Deep Graph Neural Networks
To address the scalability challenges of GNNs on large-scale graphs, deep GNNs [31, 33, 60, 61, 62] have become the dominant paradigm for graph representation learning [32, 63, 29] in recent years, owing to their powerful capability in modeling complex relational structures across various domains, such as social networks, knowledge graphs, and molecular analysis. The core idea behind GNNs is to iteratively aggregate and transform features from a node’s local neighborhood, enabling each node to capture both local and global structural information through multiple layers of message passing.
Deep GNN architectures can be broadly categorized into spectral-based and spatial-based methods. Spectral methods are grounded in graph signal processing theory and perform convolution operations in the graph spectral domain. Notably, the Graph Convolutional Network (GCN) simplifies the spectral convolution and achieves efficiency by using a localized first-order approximation, making it highly scalable for large graphs. Spatial-based methods perform aggregation directly in the graph domain, often inspired by the message passing framework. Spatial approaches have gained popularity due to their flexibility and ease of implementation for inductive learning scenarios. Additionally, Several representative deep GNN models have pushed the field forward. GraphSAGE introduces neighborhood sampling and generalizes convolution to unseen nodes, supporting inductive learning. GATs leverage attention mechanisms to adaptively weight neighbor contributions, while Graph Isomorphism Networks enhance expressive power by learning to distinguish graph structures beyond traditional message passing. In addition, more recent architectures such as PNA, GCNII, and GATv2 further improve model depth, aggregation schemes, and expressiveness.
Despite their success, deep GNNs face several significant challenges. Increasing network depth often leads to the over-smoothing problem, where node representations become indistinguishable and model performance deteriorates. Moreover, as the number of layers grows, scalability and computational complexity become major bottlenecks, especially on large-scale or dense graphs. To mitigate these issues, a variety of techniques such as residual connections, identity mapping, and attention-based neighbor selection have been proposed. In summary, deep GNN research has progressed rapidly, with continuous innovation in model architectures, aggregation mechanisms, and training paradigms. However, challenges such as over-smoothing, scalability, and effective high-order feature utilization still require further study, motivating the design of more adaptive and efficient deep GNN frameworks.
II-C Scalable Graph Neural Networks
To address the scalability challenges of GNNs on large-scale graphs, two primary research directions have emerged in scalable GNNs: sampling-based methods [64, 65, 66] and pre-computation-based techniques. While sampling reduces computation by selecting a subset of neighbors during training, it often introduces variance and information loss. In contrast, pre-computation-based approaches shift expensive operations offline, offering better stability and efficiency, and have demonstrated strong empirical performance across various benchmarks. A key idea in pre-computation is to decouple feature propagation from nonlinear transformation, allowing feature diffusion to be computed in advance without involving model parameters. SGC [20] pioneered this direction by removing nonlinearities between graph convolution layers, drastically improving scalability with minimal performance degradation. Building on this, SIGN, S2GC, GBP, and GAMLP further enhance the expressive capacity by leveraging multi-hop propagation, generalized spectral filters, or adaptive fusion, while preserving the efficiency benefits of pre-computation.
In the context of heterogeneous graphs, where multiple node and edge types introduce additional complexity, pre-computation poses new challenges. NARS [67] tackles this by sampling subsets of relations and constructing multiple subgraphs, enabling relation-specific pre-computation and reducing global information loss. SeHGNN [68] adopts a relation-wise aggregation strategy, where neighbor information is collected separately per relation within $K$ hops. This preserves finer-grained relational signals but leads to computational inefficiency as the number of relations and hops grows, particularly when $K>2$. To alleviate this trade-off, RpHGNN [69] proposes a hybrid design that combines the low information loss of the relation-wise strategy with the computational advantages of representation-wise aggregation, striking a more effective balance between accuracy and scalability in large-scale graphs.
Especially, there exist plug-in models such as RMask [41], which can be directly integrated into existing scalable GNN frameworks. These methods aim to improve both scalability and model depth by eliminating noise during feature propagation, thereby alleviating over-smoothing and reducing pre-processing overhead. By acting as flexible modules, plug-in approaches enable deeper propagation and more effective feature aggregation, significantly boosting the performance of various model-simplification GNNs across large-scale graphs.
In summary, scalable GNNs have greatly improved efficiency and applicability for large and complex graphs through sampling, pre-computation, and plug-in designs. While these methods enable faster and more scalable learning, relation-aware and adaptive aggregation further enhance flexibility for heterogeneous scenarios. However, effectively balancing scalability with information preservation, reducing sampling variance, and mitigating the over-smoothing issue remain major challenges. Continued progress relies on addressing these obstacles to achieve both scalability and high-quality representation learning.
III Preliminary
@@
Generally, we consider an undirected large-scale graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\,\mathcal{X}\}$, where $\mathcal{V}$ is the collection of nodes, $\mathcal{E}$ is the collection of edges between nodes, and $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the collection of node attributes, where $f$ is the size of the node feature vector. $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes the adjacency matrix of graph $\mathcal{G}$. $\mathbf{D}=diag(d_{1},d_{2},\ldots,d_{|\mathcal{V}|})\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the degree matrix of $\mathbf{A}$, where $d_{i}=\sum\nolimits_{{v_{j}}\in\mathcal{V}}{\mathbf{A}_{ij}}$ denotes the degree of node $v_{i}$. Given the labeled node set $\mathcal{V}_{l}$, our goal is to predict the class labels for nodes in the unlabeled set $\mathcal{V}_{u}=\mathcal{V}-\mathcal{V}_{l}$ under the supervision of $\mathcal{V}_{l}$.
IV Methodology